OpenStudy (anonymous):

HELP PLEASE A mirror with a parabolic cross section is used to collect sunlight on a pipe located at the focus of the mirror. The pipe is located 2 inches from the vertex of the mirror. Write an equation of the parabola that models the cross section of the mirror. Assume that the parabola opens upward.

OpenStudy (anonymous):

height of the parabola in inches: Take the formula to be y = ax^2 The focus of the parabola is 2 inches up, which means a/4 = 2 ==> a = 8... the equation is: y = 8x^2

OpenStudy (anonymous):

thank you , what if it was 6 inches?

OpenStudy (anonymous):

I would do the same thing right?

OpenStudy (anonymous):

Replace the 2 with a 6. Which means a/4 = 2 ==> a = 8. |||||| a/4 = 6 ==> a = 24. y = 24x^2

OpenStudy (anonymous):

why would it always be four?

OpenStudy (anonymous):

Yeah, I believe so. I think it's just the formula. I'll check very quickly for you tho.

OpenStudy (anonymous):

Noo noo , you don't have to i believe you lol I was just wondering , but thank you So much (:

OpenStudy (anonymous):

Whoops, y = ax^2 is the formula, I guess whatever the 'x' is, but at least you have the answer. Off to other post, later, welcome.

OpenStudy (confluxepic):

lol

OpenStudy (confluxepic):

@~

OpenStudy (confluxepic):

@nightwielder

OpenStudy (confluxepic):

@TheSmartOne

OpenStudy (confluxepic):

@ambi

OpenStudy (confluxepic):

@GreenCat

OpenStudy (confluxepic):

@Leader

OpenStudy (confluxepic):

@lol